# Formation of extremal Reissner-Nordström black holes: insights from numerics

### v1

@@ 1 Introduction
The third law of black hole mechanics states that it is not possible for a non-extremal black hole to become extremal in finite time in any (classical) process, no matter how idealised, involving physically reasonable matter. This was first conjectured by Bardeen, Carter and Hawking [Bardeen:1973gs], and stated more formally by Israel [Israel:1986gqz]. Recently, Kehle and Unger [Kehle:2022uvc] have proved that this conjecture is false for certain types of matter. They showed that it is possible for a massless charged scalar field to undergo spherically-symmetric gravitational collapse to form a black hole which is exactly Schwarzschild for a finite period of time but which then absorbs charge to become an exactly extremal Reissner-Nordström black hole after a finite time. They have constructed similar solutions for massless charged Vlasov matter [Kehle:2024vyt]. They have also conjectured that it should be possible to form an extremal Kerr black hole in finite time starting from vacuum initial data.
@@
The proof of [Kehle:2022uvc] is based on characteristic gluing, a method of constructing solutions to Einstein’s equations first developed by Aretakis, Czimek and Rodnianski [Aretakis:2021fel, Aretakis:2021fgz, Aretakis:2021lzi, Czimek:2022max]. Two explicitly-known spacetimes are glued along a common characteristic (i.e. null) hypersurface by choosing appropriate data on this hypersurface. At early/late time the data on this hypersurface coincides with that of the original spacetimes to be glued, with a non-trivial “gluing” region in between. The data in this region must be chosen such that the equations of motion hold on the surface and the metric and matter fields have the desired degree of smoothness, which requires that suitably many transverse derivatives of the equations of motion must also be satisfied on the surface. In general, this procedure requires fine-tuning of the data in the gluing region.
@@
In the construction of [Kehle:2022uvc], the characteristic gluing hypersurface is the event horizon. In the simplest case, it is constructed by gluing a null cone in Minkowski spacetime to the event horizon of an extremal Reissner-Nordström black hole (a more complicated 2-stage gluing is required to ensure an intermediate Schwarzschild phase). In spherical symmetry, the free data is simply the scalar field profile in the gluing region, i.e. the scalar field profile on a section of the event horizon. Kehle and Unger use an Ansatz for this profile consisting of a superposition of a finite number of disjoint bump functions, with each bump multiplied by an arbitrary parameter. They proved that there exists a choice for these parameters that achieves gluing, provided that the mass $M$ of the final black hole is large in units of the scalar field charge $\mathfrak{e}$, i.e. $\mathfrak{e}M\gg 1$. Once the characteristic data on the horizon has been constructed, general theorems can be used to prove that it arises from a spacetime describing gravitational collapse to produce a black hole that is exactly extremal Reissner-Nordström after a finite time.
@@
Although very elegant, the nature of the proof does not reveal much about the actual form of the scalar field profile on the horizon. It is also unclear just how large $\mathfrak{e}M$ has to be in order for the gluing construction to work. In this paper we will answer these questions by performing the gluing numerically, using several different Ansätze for the scalar field profile on the horizon. Some of these are of the form discussed by Kehle and Unger, i.e. disjoint bumps, but others are not. For each Ansatz we determine the minimum value of $\mathfrak{e}M$ required to achieve gluing. We find that this differs significantly for the different Ansätze. This situation is reminiscent of critical collapse [Choptuik1993, Gundlach1997, Gundlach2007], where the threshold amplitude depends sensitively on the initial profile while the critical solution itself—analogous here to the extremal black hole—remains universal.
@@
A second question that we will investigate is how the inclusion of a non-zero mass for the scalar field affects the results. The examples of Kehle and Unger involve matter with vanishing mass-to-charge ratio. Kehle and Unger argued that their results will also hold for a charged scalar field or charged Vlasov matter with a non-zero but small mass-to-charge ratio. However, if the mass-to-charge ratio is not small enough then it has been proved that examples like these cannot exist. For matter satisfying a “local mass-charge inequality”, asserting roughly that the mass-to-charge ratio of matter is greater than (or equal to) $1$ in suitable units, then it is not possible for an extremal Reissner-Nordström black hole to form in finite time in gravitational collapse [Reall:2024njy] or from a pre-existing non-extremal black hole [McSharry:2025iuz]. [^1] Similar results hold for supersymmetric asymptotically anti-de Sitter black holes [McSharry:2025iuz]. It is not known whether this result is sharp, i.e. if matter has mass-to-charge ratio strictly less than $1$ then can one construct a solution describing gravitational collapse to form an extremal Reissner-Nordström black hole in finite time? We will explore this question numerically for the case of a scalar field of mass $\mathfrak{m}$ and charge $\mathfrak{e}$. For each of our scalar field Ansätze, we will show that if $\mathfrak{m}/\mathfrak{e}$ is not too large (well below $1$ but beyond the regime of validity of the methods of Kehle and Unger) then the gluing can be performed and so there exists a solution describing gravitational collapse to form an extremal Reissner-Nordström black hole in finite time.
@@
This paper is structured as follows. In Section 2, we describe characteristic gluing, introduce the Einstein-Maxwell-charged scalar field system, and provide the corresponding governing equations in spherical symmetry. We discuss gauge freedom and how it can be exploited in our search for gluing solutions. We begin Section 3 by briefly reviewing the results of Kehle and Unger [Kehle:2022uvc]. We then describe our numerical results, starting with the lowest regularity class, $C^{0}$. We later explore $C^{1}$ gluing, discuss the $C^{2}$ case, and compare our findings of each case. In Section 4, we summarise and discuss our results and some possible future directions.
@@ 2.1 Introduction to gluing
Suppose we would like to construct a spacetime that contains a region that is (conformally) flat, and a region that is the exterior region of a stationary black hole. One approach is to use characteristic gluing: assume the two regions lie on a common null cone, and choose the metric functions and the matter fields to interpolate appropriately between the two regions along this cone. Then, by exploiting the well-posedness of the characteristic initial value problem [Luk:2011vf], one can extend the solution to a slab of finite thickness around the cone, thereby obtaining a spacetime with the desired properties.
To make this idea more precise, let us introduce some notation. [^2] We have borrowed some of the notation of [Kehle:2022uvc]. See also their excellent introduction on characteristic gluing, and references therein. Let $U$, $V$, be outgoing and ingoing null coordinates. Let $\mathfrak{R}_{1}=[0,U_{2}]\times[V_{1},0]$ and $\mathfrak{R}_{2}=[U_{1},0]\times[1,V_{2}]$ be rectangular regions, isometric to a subset of some spacetime $\mathcal{Q}_{1}$, $\mathcal{Q}_{2}$, respectively, for some $U_{1}<0<U_{2}$ and $V_{1}<0<1<V_{2}$ (see Figure 1). The two regions, by construction, touch the outgoing null cone $\mathcal{C}\equiv\{(U,V)\,|\,U=0,V\in[0,1]\}$. In each of $\mathfrak{R}_{1}$, $\mathfrak{R}_{2}$, the spacetime is known exactly. Our goal now is to determine what values the metric and matter fields should take in the other regions—in particular, on $\mathcal{C}$.
Figure 1: Setup for characteristic gluing with zero cosmological constant along the outgoing null cone $\mathcal{C}\equiv\{0\}\times[0,1]$. The metric and matter content in regions $\mathfrak{R}_{1}$ and $\mathfrak{R}_{2}$ are known and are given by $\mathcal{Q}_{1}$ and $\mathcal{Q}_{2}$, respectively. Data must be imposed on $\mathcal{C}$ such that the Einstein equations are satisfied. In the spherically-symmetric situation considered in this paper, $\mathfrak{R}_{1}$ can be extended to $r=0$, the centre of symmetry, to obtain the dark triangular region. The rectangle $\mathfrak{R}_{2}$ can be extended to get a complete future null infinity $\mathcal{I}^{+}$, as well as a segment of past null infinity $\mathcal{I}^{-}$, and spacelike ($i^{0}$) and future timelike ($i^{+}$) infinity. Once a valid solution to the characteristic gluing problem is found, the well-posedness of the characteristic initial value problem [Luk:2011vf] can be exploited to obtain a local extension of spacetime. To extend this local solution to $r=0$, one can use Cauchy stability [Kehle:2022uvc]. A Cauchy surface is shown in red.
@@
Once $\mathfrak{R}_{1}$ and $\mathfrak{R}_{2}$ have been appropriately glued, the null cone $\mathcal{C}$ can be extended into a full $4$-dimensional region of spacetime, thanks to a result of Luk [Luk:2011vf]. Taking first the two null hypersurfaces $\{(U,V)\,|\,U\geq 0,V=0\}$ and $\{(U,V)\,|\,U=0,V\geq 0\}$, one can locally extend the spacetime to the north. Next, taking the hypersurfaces $\{(U,V)\,|\,U\leq 0,V=1\}$ and $\{(U,V)\,|\,U=0,V\leq 1\}$, one can locally extend the spacetime to the south. Extending this segment to the centre, $r=0$, is non-trivial, but can be achieved by exploiting Cauchy stability [Kehle:2022uvc]. Note that, in spherical symmetry, the extended spacetime solution loses a derivative at the centre compared with the characteristic data [Kehle:2022uvc]. Thus, if we wish to obtain a spacetime of regularity $C^{k}$, we require $C^{k+1}$ characteristic data.
@@
In this paper, we will focus on determining the appropriate data to impose on $\mathcal{C}$, but will not discuss the extension to a full spacetime in any further detail, as this is already well explained in Section 5 of [Kehle:2022uvc].
@@ 2.2 Einstein-Maxwell-Charged Scalar Field System
Let us now make the discussion more concrete by introducing the system we will study: the Einstein-Maxwell-charged scalar field system. The equations of motion read [Kehle:2022uvc, Rossetti:2023mrx]
@@ 2.3 Gauge Freedom
There is some residual gauge freedom in our Ansatz (3) which we can use to simplify the problem [Kehle:2022uvc]. Firstly, we can rescale $U\to f(U)$, $V\to g(V)$, so that the $dUdV$ term in the metric becomes
@@
where $r(U,V)=(V-U)/2\in[0,\infty)$. Thus we can read off $\Omega\equiv 1$, $\partial_{U}r=-1/2$, $\partial_{V}r=1/2$, $\mathcal{A}_{U}=0$, $\Phi\equiv 0$, and of course the Hawking mass and the charge both vanish. After performing a gauge transformation of the type described above, we can bring $\partial_{U}r$ to any negative value. Since $m$ is gauge-invariant, $\partial_{V}r$ must transform such that $m$ remains zero. Similarly, we can transform $\mathcal{A}_{U}$ to any desired value, but $Q$ must remain zero. We therefore obtain the following result [Kehle:2022uvc]:
@@
A lapse-normalised sphere on the horizon of a Reissner-Nordström black hole has $\Omega\equiv 1$, $r=r_{+}$, $\partial_{U}r=-1/2$, $\partial_{V}r=0$, $Q=\mathfrak{q}M$, $\Phi\equiv 0$. Hence we have [Kehle:2022uvc]:
@@ 2.4 Scalar field Ansätze
In practice, as in [Kehle:2022uvc], we begin by first solving the Raychaudhuri equation (5a) backwards from $V=1$ to $V=0$ (using the initial conditions $r(1)=r_{+}$, $\partial_{V}r(1)=0$). Then, we evolve the remaining equations, (5c)–(5h), forwards in time. The initial condition for $\partial_{U}r(0)$ can be derived from the requirement that the Hawking mass be zero, giving
@@
Without loss of generality, let us assume that $\mathfrak{e}M>0$. Suppose that we wish to obtain a black hole with $\mathfrak{q}>0$. Following [Kehle:2022uvc], we choose an Ansatz for $\Phi$ of the form
@@
We have mainly experimented with three different choices for the basis functions $\chi_{j}$. The first two, inspired by [Kehle:2022uvc], consist of compactly-supported bump functions with disjoint support. More specifically, in the region they are non-zero, these functions are
@@ 3.1 General approach
Following [Kehle:2022uvc], we will write the set of parameters $\alpha_{j}$ as a vector $\alpha=(\alpha_{1},\dots,\alpha_{2k+1})\in\mathbb{R}^{2k+1}$ and denote its norm by $\beta\equiv|\alpha|$. We will also denote the unit-norm vector pointing in the same direction as $\alpha$ by $\hat{\alpha}\in S^{2k}$, so that we have $\alpha=\beta\hat{\alpha}$. $C^{k}$ gluing solutions are constructed as follows.
@@ 3.2.1 Candidate solutions
In $C^{0}$ gluing, the Ansatz for the scalar field on $\mathcal{C}$ involves a single free parameter $\alpha_{1}$. The first step is to determine the restricted parameter space defined above. Numerically we find that this is given by $|\alpha_{1}|<\alpha_{\rm max}$ where $\alpha_{\rm max}=0.8391$, $0.4225$ and $0.6141$ for the even, odd and polynomial Ansätze, respectively. For each $\alpha_{1}$ in the restricted parameter space we can calculate $I(\alpha)$; the result is given in the first panel of Figure 3. We find that $I_{\rm max}\equiv\sup_{\alpha}I$ takes the values $I_{\rm max}=0.0627$, $0.0186$ and $0.0934$ for the even, odd and polynomial Ansätze, respectively. In $C^{0}$ gluing we do not need to match derivatives of the scalar field and so the only equation to be solved by a candidate solution is (21). Note that $\mathfrak{m}$ does not enter this equation so $\mathfrak{m}$ does not affect candidate solutions of $C^{0}$ gluing. This equation admits a solution if, and only if,

### v2

@@ 1 Introduction
The third law of black hole mechanics states that it is not possible for a non-extremal black hole to become extremal in finite time in any (classical) process, no matter how idealised, involving physically reasonable matter. This was first conjectured by Bardeen, Carter and Hawking [4], and stated more formally by Israel [11]. Recently, Kehle and Unger [12] have proved that this conjecture is false for certain types of matter. They showed that it is possible for a massless charged scalar field to undergo spherically-symmetric gravitational collapse to form a black hole which is exactly Schwarzschild for a finite period of time but which then absorbs charge to become an exactly extremal Reissner-Nordström black hole after a finite time. They have constructed similar solutions for massless charged Vlasov matter [13]. They have also conjectured that it should be possible to form an extremal Kerr black hole in finite time starting from vacuum initial data.
@@
The proof of [12] is based on characteristic gluing, a method of constructing solutions to Einstein’s equations first developed by Aretakis, Czimek and Rodnianski [2, 1, 3, 7]. Two explicitly-known spacetimes are glued along a common characteristic (i.e. null) hypersurface by choosing appropriate data on this hypersurface. At early/late time the data on this hypersurface coincides with that of the original spacetimes to be glued, with a non-trivial “gluing” region in between. The data in this region must be chosen such that the equations of motion hold on the surface and the metric and matter fields have the desired degree of smoothness, which requires that suitably many transverse derivatives of the equations of motion must also be satisfied on the surface. In general, this procedure requires fine-tuning of the data in the gluing region.
@@
In the construction of [12], the characteristic gluing hypersurface is the event horizon. In the simplest case, it is constructed by gluing a null cone in Minkowski spacetime to the event horizon of an extremal Reissner-Nordström black hole (a more complicated 2-stage gluing is required to ensure an intermediate Schwarzschild phase). In spherical symmetry, the free data is simply the scalar field profile in the gluing region, i.e. the scalar field profile on a section of the event horizon. Kehle and Unger use an Ansatz for this profile consisting of a superposition of a finite number of disjoint bump functions, with each bump multiplied by an arbitrary parameter. They proved that there exists a choice for these parameters that achieves gluing, provided that the mass $M$ of the final black hole is large in units of the scalar field charge $\mathfrak{e}$, i.e. $\mathfrak{e}M\gg 1$. Once the characteristic data on the horizon has been constructed, general theorems can be used to prove that it arises from a spacetime describing gravitational collapse to produce a black hole that is exactly extremal Reissner-Nordström after a finite time.
@@
Although very elegant, the nature of the proof does not reveal much about the actual form of the scalar field profile on the horizon. It is also unclear just how large $\mathfrak{e}M$ has to be in order for the gluing construction to work. In this paper we will answer these questions by performing the gluing numerically, using several different Ansätze for the scalar field profile on the horizon. Some of these are of the form discussed by Kehle and Unger, i.e. disjoint bumps, but others are not. For each Ansatz we determine the minimum value of $\mathfrak{e}M$ required to achieve gluing. We find that this differs significantly for the different Ansätze. This situation is reminiscent of critical collapse [6, 9, 10], where the threshold amplitude depends sensitively on the initial profile while the critical solution itself—analogous here to the extremal black hole—remains universal.
@@
A second question that we will investigate is how the inclusion of a non-zero mass for the scalar field affects the results. The examples of Kehle and Unger involve matter with vanishing mass-to-charge ratio. Kehle and Unger argued that their results will also hold for a charged scalar field or charged Vlasov matter with a non-zero but small mass-to-charge ratio. However, if the mass-to-charge ratio is not small enough then it has been proved that examples like these cannot exist. For matter satisfying a “local mass-charge inequality”, asserting roughly that the mass-to-charge ratio of matter is greater than (or equal to) $1$ in suitable units, then it is not possible for an extremal Reissner-Nordström black hole to form in finite time in gravitational collapse [17] or from a pre-existing non-extremal black hole [16]. [^1] Similar results hold for supersymmetric asymptotically anti-de Sitter black holes [16]. It is not known whether this result is sharp, i.e. if matter has mass-to-charge ratio strictly less than $1$ then can one construct a solution describing gravitational collapse to form an extremal Reissner-Nordström black hole in finite time? We will explore this question numerically for the case of a scalar field of mass $\mathfrak{m}$ and charge $\mathfrak{e}$. For each of our scalar field Ansätze, we will show that if $\mathfrak{m}/\mathfrak{e}$ is not too large (well below $1$ but beyond the regime of validity of the methods of Kehle and Unger) then the gluing can be performed and so there exists a solution describing gravitational collapse to form an extremal Reissner-Nordström black hole in finite time.
@@
This paper is structured as follows. In Section 2, we describe characteristic gluing, introduce the Einstein-Maxwell-charged scalar field system, and provide the corresponding governing equations in spherical symmetry. We discuss gauge freedom and how it can be exploited in our search for gluing solutions. We begin Section 3 by briefly reviewing the results of Kehle and Unger [12]. We then describe our numerical results, starting with the lowest regularity class, $C^{0}$. We later explore $C^{1}$ gluing, discuss the $C^{2}$ case, and compare our findings of each case. In Section 4, we summarise and discuss our results and some possible future directions.
@@ 2.1 Introduction to gluing
Suppose we would like to construct a spacetime that contains a region that is (conformally) flat, and a region that is the exterior region of a stationary black hole. One approach is to use characteristic gluing: assume the two regions lie on a common null cone, and choose the metric functions and the matter fields to interpolate appropriately between the two regions along this cone. Then, by exploiting the well-posedness of the characteristic initial value problem [14], one can extend the solution to a slab of finite thickness around the cone, thereby obtaining a spacetime with the desired properties.
To make this idea more precise, let us introduce some notation. [^2] We have borrowed some of the notation of [12]. See also their excellent introduction on characteristic gluing, and references therein. Let $U$, $V$, be outgoing and ingoing null coordinates. Let $\mathfrak{R}_{1}=[0,U_{2}]\times[V_{1},0]$ and $\mathfrak{R}_{2}=[U_{1},0]\times[1,V_{2}]$ be rectangular regions, isometric to a subset of some spacetime $\mathcal{Q}_{1}$, $\mathcal{Q}_{2}$, respectively, for some $U_{1}<0<U_{2}$ and $V_{1}<0<1<V_{2}$ (see Figure 1). The two regions, by construction, touch the outgoing null cone $\mathcal{C}\equiv\{(U,V)\,|\,U=0,V\in[0,1]\}$. In each of $\mathfrak{R}_{1}$, $\mathfrak{R}_{2}$, the spacetime is known exactly. Our goal now is to determine what values the metric and matter fields should take in the other regions—in particular, on $\mathcal{C}$.
Figure 1: Setup for characteristic gluing with zero cosmological constant along the outgoing null cone $\mathcal{C}\equiv\{0\}\times[0,1]$. The metric and matter content in regions $\mathfrak{R}_{1}$ and $\mathfrak{R}_{2}$ are known and are given by $\mathcal{Q}_{1}$ and $\mathcal{Q}_{2}$, respectively. Data must be imposed on $\mathcal{C}$ such that the Einstein equations are satisfied. In the spherically-symmetric situation considered in this paper, $\mathfrak{R}_{1}$ can be extended to $r=0$, the centre of symmetry, to obtain the dark triangular region. The rectangle $\mathfrak{R}_{2}$ can be extended to get a complete future null infinity $\mathcal{I}^{+}$, as well as a segment of past null infinity $\mathcal{I}^{-}$, and spacelike ($i^{0}$) and future timelike ($i^{+}$) infinity. Once a valid solution to the characteristic gluing problem is found, the well-posedness of the characteristic initial value problem [14] can be exploited to obtain a local extension of spacetime. To extend this local solution to $r=0$, one can use Cauchy stability [12]. A Cauchy surface is shown in red.
@@
Once $\mathfrak{R}_{1}$ and $\mathfrak{R}_{2}$ have been appropriately glued, the null cone $\mathcal{C}$ can be extended into a full $4$-dimensional region of spacetime, thanks to a result of Luk [14]. Taking first the two null hypersurfaces $\{(U,V)\,|\,U\geq 0,V=0\}$ and $\{(U,V)\,|\,U=0,V\geq 0\}$, one can locally extend the spacetime to the north. Next, taking the hypersurfaces $\{(U,V)\,|\,U\leq 0,V=1\}$ and $\{(U,V)\,|\,U=0,V\leq 1\}$, one can locally extend the spacetime to the south. Extending this segment to the centre, $r=0$, is non-trivial, but can be achieved by exploiting Cauchy stability [12]. Note that, in spherical symmetry, the extended spacetime solution loses a derivative at the centre compared with the characteristic data [12]. Thus, if we wish to obtain a spacetime of regularity $C^{k}$, we require $C^{k+1}$ characteristic data.
@@
In this paper, we will focus on determining the appropriate data to impose on $\mathcal{C}$, but will not discuss the extension to a full spacetime in any further detail, as this is already well explained in Section 5 of [12].
@@ 2.2 Einstein-Maxwell-Charged Scalar Field System
Let us now make the discussion more concrete by introducing the system we will study: the Einstein-Maxwell-charged scalar field system. The equations of motion read [12, 18]
@@ 2.3 Gauge Freedom
There is some residual gauge freedom in our Ansatz (3) which we can use to simplify the problem [12]. Firstly, we can rescale $U\to f(U)$, $V\to g(V)$, so that the $dUdV$ term in the metric becomes
@@
where $r(U,V)=(V-U)/2\in[0,\infty)$. Thus we can read off $\Omega\equiv 1$, $\partial_{U}r=-1/2$, $\partial_{V}r=1/2$, $\mathcal{A}_{U}=0$, $\Phi\equiv 0$, and of course the Hawking mass and the charge both vanish. After performing a gauge transformation of the type described above, we can bring $\partial_{U}r$ to any negative value. Since $m$ is gauge-invariant, $\partial_{V}r$ must transform such that $m$ remains zero. Similarly, we can transform $\mathcal{A}_{U}$ to any desired value, but $Q$ must remain zero. We therefore obtain the following result [12]:
@@
A lapse-normalised sphere on the horizon of a Reissner-Nordström black hole has $\Omega\equiv 1$, $r=r_{+}$, $\partial_{U}r=-1/2$, $\partial_{V}r=0$, $Q=\mathfrak{q}M$, $\Phi\equiv 0$. Hence we have [12]:
@@ 2.4 Scalar field Ansätze
In practice, as in [12], we begin by first solving the Raychaudhuri equation (5a) backwards from $V=1$ to $V=0$ (using the initial conditions $r(1)=r_{+}$, $\partial_{V}r(1)=0$). Then, we evolve the remaining equations, (5c)–(5h), forwards in time. The initial condition for $\partial_{U}r(0)$ can be derived from the requirement that the Hawking mass be zero, giving
@@
Without loss of generality, let us assume that $\mathfrak{e}M>0$. Suppose that we wish to obtain a black hole with $\mathfrak{q}>0$. Following [12], we choose an Ansatz for $\Phi$ of the form
@@
We have mainly experimented with three different choices for the basis functions $\chi_{j}$. The first two, inspired by [12], consist of compactly-supported bump functions with disjoint support. More specifically, in the region they are non-zero, these functions are
@@ 3.1 General approach
Following [12], we will write the set of parameters $\alpha_{j}$ as a vector $\alpha=(\alpha_{1},\dots,\alpha_{2k+1})\in\mathbb{R}^{2k+1}$ and denote its norm by $\beta\equiv|\alpha|$. We will also denote the unit-norm vector pointing in the same direction as $\alpha$ by $\hat{\alpha}\in S^{2k}$, so that we have $\alpha=\beta\hat{\alpha}$. $C^{k}$ gluing solutions are constructed as follows.
@@ 3.2.1 Candidate solutions
In $C^{0}$ gluing, the Ansatz for the scalar field on $\mathcal{C}$ involves a single free parameter $\alpha_{1}$. The first step is to determine the restricted parameter space defined above. Numerically we find that this is given by $|\alpha_{1}|<\alpha_{\rm max}$ where $\alpha_{\rm max}=0.5933$, $0.2987$ and $0.4343$ for the even, odd and polynomial Ansätze, respectively. For each $\alpha_{1}$ in the restricted parameter space we can calculate $I(\alpha)$; the result is given in the first panel of Figure 3. We find that $I_{\rm max}\equiv\sup_{\alpha}I$ takes the values $I_{\rm max}=0.0627$, $0.0186$ and $0.0934$ for the even, odd and polynomial Ansätze, respectively. In $C^{0}$ gluing we do not need to match derivatives of the scalar field and so the only equation to be solved by a candidate solution is (21). Note that $\mathfrak{m}$ does not enter this equation so $\mathfrak{m}$ does not affect candidate solutions of $C^{0}$ gluing. This equation admits a solution if, and only if,